OpenStudy (anonymous):

Two Storage Sheds Are To Have The Same Area. One Is Square And One Is Rectangular. The Rectangular Shed Is 3 Meters Wide And 12 Meters Long. How Large Is One Side Of The Square Shed In Meters

hartnn (hartnn):

area of rectangle = ?

OpenStudy (anonymous):

They Have The Same Area

hartnn (hartnn):

u don't know the formula of area of a rectangle ??

OpenStudy (anonymous):

Area is Same ie l*b = a^2

OpenStudy (anonymous):

12*3 = a^2 a =6m

OpenStudy (anonymous):

so...SIde of Square is 6m

OpenStudy (anonymous):

Thanks @Yahoo! And @hartnn

OpenStudy (anonymous):

Welcome...

hartnn (hartnn):

don't thank me, i didn't help.

hartnn (hartnn):

the great yahoo only helped